# Motion-Aware Generative Frame Interpolation

### v1

@@ 5 Experiment
GI [33] 91.81 91.28 91.70 94.16 93.86 96.43 97.51 98.00 47.44 51.13 57.51 60.84 79.97 81.97
TRF [3] 89.25 89.25 92.26 94.01 93.48 95.77 96.77 97.18 47.94 50.12 54.38 58.40 79.01 80.79
DynamiCrafter [35] 89.06 91.94 92.97 96.39 48.47 58.24 79.51
ToonCrafter [34] 91.78 95.42 95.63 96.75 51.44 64.37 82.57
MoG (ours) 92.65 92.91 94.34 96.11 94.69 96.58 97.90 96.94 49.84 52.43 59.23 64.89 81.44 83.31
Table 1: Quantitative comparison on video quality. Bold text indicates the best results in generative frame interpolation.
[…]
Real Anime Real Anime Real Anime Real Anime Real Anime Real Anime
      Flow-based VFI models
RIFE [10] 18.13 20.27 0.6094 0.7817 0.3745 0.3407 898.38 647.49 65.04 66.16 0.8228 0.8643
EMA-VFI [41] 18.11 20.45 0.6118 0.7843 0.3763 0.3701 808.54 544.48 60.28 55.84 0.8363 0.8895
      Generative VFI models
GI [33] 15.95 18.04 0.5271 0.6971 0.3384 0.2891 521.00 449.31 36.06 46.18 0.8703 0.8710
TRF [3] 15.43 16.49 0.5132 0.6744 0.3920 0.3470 624.63 481.02 42.48 53.95 0.8491 0.8731
DynamiCrafter [35] 16.05 0.5225 0.3380 562.34 42.16 0.8634
ToonCrafter [34] 18.01 0.7182 0.2944 425.71 40.63 0.9203
MoG (ours) 17.82 19.44 0.5898 0.7434 0.2716 0.2615 401.49 351.41 31.26 33.73 0.9083 0.9320
Table 2: Quantitative comparison on generation fidelity to the ground truth.
@@ 5.2 VFIBench
A well-interpolated video should not only be of high quality inherently but also maintain fidelity to the ground truth. For the former, we utilize six metrics from VBench [11]: subject consistency, background consistency, temporal flickering, motion smoothness, aesthetic quality, and imaging quality. These metrics collectively assess the intrinsic quality of the video. For the latter, we employ six widely adopted metrics: PSNR, SSIM, LPIPS [43], FID [7], and the CLIP similarity score [24] for image-level comparison, and FVD [29, 30] for video-level comparison.
@@
5.3 Comparative Analysis
We evaluate our MoG by benchmarking it against state-of-the-art methods across two categories: flow-based interpolation methods, specifically RIFE [10] and EMA-VFI [41], and generative interpolation methods, including GI [33], TRF [3], DynamiCrafter [35] and ToonCrafter [34]. Note that DynamiCrafter and ToonCrafter are tailored for real-world and cartoon animation data, respectively, and their performance is reported separately for each data type.
@@ 5.3 Comparative Analysis
As shown in Tab. 1 and Tab. 2, compared to others generative VFI methods, MoG exhibits significant improvements in video quality and fidelity to ground truth, particularly in consistency-related metrics. This demonstrates that the introduced motion guidance indeed enables the generative model to better understand the correspondence between input frames. Compared to flow-based VFI models, our approach also demonstrates notable enhancements across most metrics; however, it lags in PSNR, SSIM, Temporal Flickering, and Motion Smoothness. We argue that this discrepancy arises because flow-based VFI often produces blurry results in complex motion scenarios (as illustrated in Fig. 1), which can inflate these metrics while compromising actual visual quality [43].
Qualitative results.
We present qualitative comparisons with three generative VFI methods in Fig. 4. Lacking explicit motion guidance, these methods struggle to accurately infer and understanding the correspondences between input frames, resulting in inconsistent content and unstable motion. In contrast, MoG achieves superior motion and visual quality in complex scenarios. More comparisons are available on the provided website in supplementary materials.
User study.
5.4 Ablation Study
@@
Effectiveness of motion guidance.
As a key contribution of our work, we thoroughly validate the efficacy of the proposed motion guidance. In Fig. 5, (a) displays the results from the DynamiCrafter, which suffers from significant motion instability and content inconsistency. (b) presents outcomes warped by the intermediate flow from [41]. Although noticeable blurring is evident, the warped intermediate frames offer valuable approximations of both motion and appearance. (c) shows results obtained through direct guidance integration using our method in a train-free manner, demonstrating substantial improvements in motion stabilization. However, this direct integration, lacking training, leads to alterations in feature distribution that affect the color and details of generated frames. Finally, (d) showcases the results of the MoG, which both incorporates guidance and fine-tunes the spatial layers. Our method effectively leverages this guidance, producing realistic motion trajectories and maintaining consistent content. More comparisons can be found in the supplementary materials.
Figure 5: Ablation on motion guidance integration.
Choice of motion guidance.
To select the appropriate motion guidance, we evaluate various approaches in Sec. 4.1 and compare their results with those obtained without motion guidance (only fine-tuning), as shown in Tab. 4(a). As discussed in Sec. 4.1, linear interpolation offers only marginal improvements in video quality and fidelity, whereas the intermediate flow from flow-based VFI achieves superior performance due to its task-oriented training and explicit occlusion modeling.
Different levels of guidance injection.
@@
We introduce motion guidance at both the latent and feature levels. To validate the effectiveness of each level, we compare the performance of models with motion guidance introduced at only one level, or without motion guidance, as in Tab. 4(d). The results demonstrate that the injection of either level all significantly enhances video quality and fidelity metrics. Specifically, latent-level injection is more beneficial for background consistency, while feature-level injection improves subject consistency. The best performance is achieved when both levels are employed, allowing the generative model to leverage motion guidance at different granularities to facilitate the generation of intermediate frames.
Design details of feature-level injection.
@@
We investigate the details of feature-level injection from two perspectives: how and where to merge motion guidance. Regarding how to merge guidance, we also devise methods that first concatenate the warped intermediate features with the original features, followed by learnable modules such as Transformer blocks, convolutions, or linear layers. Surprisingly, as shown in Tab. 4(c), a simple averaging yielded the best performance, possibly due to the substantial data requirements for the learnable modules to achieve strong generalization. In terms of where to merge motion guidance, we experiment with three configurations: injecting into all blocks (All), exclusively into expansive blocks (Decoder-only), and solely into contracting blocks (Encoder-only). As illustrated in Tab. 4(b), the Encoder-only achieves the best performance, while Decoder-only results in the poorest performance. This discrepancy may stem from the fact that modifications to the decoder can disrupt the powerful video generation capabilities from pre-training.
@@ 6 Conclusion
In this work, we have presented a novel generative frame interpolation framework, MoG, which enhances the model’s motion awareness by explicitly incorporating additional motion guidance. We first reveal that the intermediate flow in flow-based VFI can serve as a suitable motion guidance. Subsequently, we propose a simple yet effective strategy to inject this guidance into the generative network at both the latent and feature levels, significantly reducing the challenge of inferring the motion of input frames. Through extensive experiments in both real-world and animated scenes, we demonstrate that MoG achieves substantial improvements in video quality and fidelity. User studies further indicate that MoG exhibits superior overall visual quality.
@@
Figure 6: Additional ablation experiments on motion guidance integration.
@@ A. More Visual Comparisons
To further demonstrate the improvement of our method, we provide video comparisons of different methods on the provided website. We also provide additional qualitative comparisons in Fig. 8.
B. Effectiveness of Introducing Motion Guidance
As shown in Fig. 6, we provide more examples similar to Fig. 5 to demonstrate the effectiveness of introducing motion guidance. By explicitly incorporating a rough estimate of the motion between input frames, our method effectively stabilizes the generated motion and reduces the inconsistency in content between the generated video and the input frames.
C. Discussion on Ineffective Motion Guidance
Our method significantly enhances video generation quality primarily by leveraging the coarse intermediate frame estimates provided by motion guidance. To present a comprehensive analysis, it is essential to demonstrate our method’s performance in scenarios where motion guidance is ineffective. As illustrated in the two examples in Fig. 7, when the motion in the input frames is overly complex, the intermediate flow fails to deliver effective motion guidance. Notably, despite the ineffectiveness of motion guidance in these cases, our approach still yields high-quality results. This indicates that our method, MoG, does not rely entirely on motion guidance for generating identical motion. Instead, it dynamically integrates the effective information from motion guidance to assist in the inference of real motion. This aligns perfectly with our original intent.
D. Limitations and Future Work
Despite MoG has achieved non-trivial improvement in generation quality across various scenes, there still several limitations warrant further exploration. Firstly, our approach is built upon the U-Net architecture of the DynamiCrafter model. However, the video generation capabilities of DynamiCrafter has lagged behind recently DiT-based [22] video generation models [40, 4], which constrains our performance ceiling. Investigating our method within the new framework presents a promising avenue for future work. Secondly, our approach relies on a pre-trained flow-based VFI model, meaning that the quality of its outputs may impact the effectiveness of motion guidance. For instance, flow-based VFI struggles to establish correspondences for extremely large motion [15]. Enhancing the generalizability of flow-based VFI across diverse scenes will also benefit our method moving forward.
Figure 7: Instances of ineffective guidance.

### v2

@@ 5 Experiment
LDMVFI [3] 17.17 18.39 0.5953 0.7175 0.3081 0.2860 41.47 46.18 0.8703 0.8710 479.63 435.17 78.91 80.21
GI [37] 15.95 18.04 0.5271 0.6971 0.3384 0.2891 36.06 46.18 0.8703 0.8710 521.00 449.31 79.97 81.97
TRF [4] 15.43 16.49 0.5132 0.6744 0.3920 0.3470 42.48 53.95 0.8491 0.8731 624.63 481.02 79.01 80.79
DynamiCrafter [39] 16.05 0.5225 0.3380 42.16 0.8634 562.34 79.51
ToonCrafter [38] 18.01 0.7182 0.2944 40.63 0.9203 425.71 82.57
MoG (ours) 17.82 19.44 0.5898 0.7434 0.2716 0.2615 31.26 33.73 0.9083 0.9320 401.49 351.41 81.44 83.31
Table 1: Quantitative comparison on VFIBench.
@@ 5.2 VFIBench
A well-interpolated video should not only be of high quality inherently but also maintain fidelity to the ground truth. For the former, we adopt six metrics from VBench [10]: subject consistency, background consistency, temporal flickering, motion smoothness, aesthetic quality, and imaging quality. The average performance across all metrics is reported as VBench in Tab. 1. These metrics collectively assess the intrinsic quality. For the latter, we employ six widely adopted metrics: PSNR, SSIM, LPIPS [48], FID [7], and the CLIP similarity score [27] for image-level comparison, and FVD [32, 33] for video-level comparison.
@@
5.3 Main Results and Analysis
We evaluate our MoG by benchmarking it against state-of-the-art methods across two categories: flow-based interpolation methods, specifically RIFE [9] and EMA-VFI [45], and generative interpolation methods, including LDMVFI [3], GI [37], TRF [4], DynamiCrafter [39] and ToonCrafter [38]. To conduct a more equitable comparison, we retrain the flow-based methods and LDMVFI [3] on our dataset. Note that DynamiCrafter and ToonCrafter are tailored for real-world and animation, respectively. Hence their performance is reported separately for each domain.
@@ 5.3 Main Results and Analysis
As shown in Tab. 1, compared to others generative VFI methods, MoG exhibits significant improvements in video quality and fidelity to ground truth. This indicates that our method can effectively utilize the intermediate flow to generate smooth motion, and meanwhile, it has successfully reduced the erroneous information within it. Compared to flow-based VFI models, our approach also demonstrates notable enhancements across most metrics; however, it lags in PSNR. We argue that this discrepancy arises because flow-based VFI often produces blurry results in complex motion scenarios (as illustrated in Fig. 1), which can inflate these metrics while compromising actual visual quality [48]. We also conduct a comparison on the publicly available dataset Davis-7 [12]. As shown in Tab. 2, MoG achieves best performance across all metrics.
  Davis-7 SSIM ($\uparrow$) LPIPS ($\downarrow$) FID ($\downarrow$)
  LDMVFI [3] 0.4175 0.2765 56.28
VIDIM [12] 0.4221 0.2986 53.38
DynamiCrafter [39] 0.4785 0.3752 75.06
MoG (ours) 0.5978 0.2641 51.94
Table 2: Comparsion on Davis-7 [12].
@@
  GI [37] TRF [4] DynamiCrafter [39] MoG (ours)
  Runtime (s) 385.19 141.41 33.42 34.08
VBench ($\uparrow$) 79.97 79.01 79.51 81.44
Table 4: Comparsion on computational efficiency.
Qualitative results.
We present qualitative comparisons with three generative VFI methods in Fig. 4. Lacking explicit motion guidance, these methods struggle to accurately infer and understanding the correspondences between input frames, resulting in inconsistent content and unstable motion. In contrast, MoG achieves superior motion and visual quality in complex scenarios. More comparisons are available in supplementary materials.
User study.
Computational efficiency .
We further compare the computational efficiency of different generative frame interpolation methods. As shown in Tab. 4, the testing resolution is unified as $16\times 512\times 320$. Compared with TRF and GI, MoG not only exhibits superior performance but also significantly improves computational efficiency. When compared with our baseline model, DynamicCrafter, MoG have substantially enhanced the quality of the generated videos while only slightly increasing the computational load.
5.4 Ablation Study
@@
Figure 5: An example of manually altering intermediate flow.
Effectiveness of intermediate flows.
To validate that intermediate flows can indeed impose motion constraints on generative models, we attempt to manually modify the intermediate flows to control the motion. As depicted in Fig. 5, given the same start and end frames, MoG would generate a nearly static video by default. When we manually alter the intermediate flows of the ball to move vertically, the generated frames change accordingly.
In addition, we explore two other potential forms of motion guidance. One is the linear interpolation of the input frames, and the other is the flow from a pre-trained optical flow estimator [40]. As shown in Fig. 6, the intermediate flows exhibit smoother motion and fewer errors. We attribute this to the fact that intermediate flows are task-oriented flows [41], which are more suitable for the frame interpolation. Moreover, the predicted occlusion masks in Eq. 1 can also enhance the accuracy of the guidance. The performance advantages presented in Tab. 5(a) further verify this conclusion.
Figure 6: Visualizations on different methods as guidance.
Dual guidance injection.
@@
We introduce motion guidance at both the latent and feature levels. To validate the effectiveness of each level, we compare the performance with motion guidance introduced at only one level, or without motion guidance, as in Tab. 5(d). The results demonstrate that the injection of either level all significantly enhances video quality and fidelity metrics. Specifically, latent-level injection is more beneficial for background consistency, while feature-level injection improves subject consistency. The best performance is achieved when both levels are employed, allowing the generative model to leverage motion guidance at different granularities.
@@
Regarding the design of feature-level injection, we also attempt methods that first concatenate the warped intermediate features with the original features, followed by learnable modules such as Transformer blocks [34], convolutions, or linear layers. Surprisingly, as shown in Tab. 5(c), a simple averaging yielded the best performance, possibly due to the substantial data requirements for the learnable modules to achieve strong generalization.
Guidance correction.
We conduct comprehensive ablation experiments on designs proposed in Sec. 4.3. First, regarding the integration location of motion guidance, we evaluate three configurations: guidance injection into all blocks (All), exclusive injection into decoder blocks (Decoder-only), and sole injection into encoder blocks (Encoder-only). As demonstrated in Tab. 5(b), Encoder-only achieves the optimal performance, while Decoder-only yields the lowest performance. This discrepancy aligns with our hypothesis that maintaining decoder guidance-free allows the model to effectively rectify flow-constrained information from the encoder, ensuring the generated videos adhere more closely to pre-trained distributions.
Furthermore, as illustrated in Fig. 7, we validate the effectiveness of selective parameter fine-tuning. Results show that while intermediate flow guidance can improve temporal smoothness, it still leads to artifacts in complex regions. Selective fine-tuning substantially enhances the generated video quality by enabling dynamic guidance adaptation in spatial layers, while preserving the motion generation capabilities through frozen temporal layers.
Figure 7: Visualizations on selective parameter fine-tuning.
@@ 6 Conclusion
In this work, we present a novel generative frame interpolation framework, MoG, which simultaneously enhances the motion smoothness by intermediate flows and adaptively correcting flow errors through generative refinement. We first propose dual guidance injection to introduce flow constraints into the generative models at both the latent and feature levels. Then we conduct encoder-only guidance injection and selective parameters fine-tuning for guidance correction. Through extensive experiments in both real-world and animated scenes, we demonstrate that MoG achieves substantial improvements in video quality and fidelity.
@@ A. More Visual Comparisons
To further demonstrate the improvement of our method, we also provide additional qualitative comparisons in Fig. 8.
B. Limitations and Future Work
Despite MoG has achieved non-trivial improvement in generation quality across various scenes, there still several limitations warrant further exploration. Firstly, our approach is built upon the U-Net architecture of the DynamiCrafter model. However, the video generation capabilities of DynamiCrafter has lagged behind recently DiT-based [25] video generation models [44, 5], which constrains our performance ceiling. Unfortunately, we currently lack the necessary resources and data to work with the latest models. Investigating our method within the new models presents a promising direction for future work. Secondly, MoG relies on the flow-based VFI model, meaning that the quality of its outputs may impact the effectiveness of motion guidance. Enhancing the generalizability of flow-based VFI across diverse scenes will also benefit our method moving forward.